Lancaster Lumber buys $8 million of materials (net of discounts) on terms of 3/5, net 50; and it currently pays on the 5th day and takes discounts. Lancaster plans to expand, which will require additional financing. Assume 365 days in year for your calculations.

  1. If Lancaster decides to forgo discounts, how much additional credit could it obtain? Write out your answer completely. For example, 5 million should be entered as 5,000,000. Do not round intermediate calculations. Round your answer to the nearest cent.
    $  

  2. What would be the nominal cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
      %

  3. What would be the effective cost of that credit? Do not round intermediate calculations. Round your answer to two decimal places.
      %

  4. If the company could get the funds from a bank at a rate of 8%, interest paid monthly, based on a 365-day year, what would be the effective cost of the bank loan? Do not round intermediate calculations. Round your answer to two decimal places.
      %

  5. Should Lancaster use bank debt or additional trade credit?
    -Select-Bank debtAdditional trade credit
Answer #1
  1. If Lancaster decides to forgo discounts, how much additional credit could it obtain?

Answer: If Lancaster takes the discount, then it will receives a discount of 3%, which amounts to a saving of 8,000,000*3% = 240,000. Therefore, if Lancaster forgoes the discount, then it can have additional credit of 240,000, and has to pay within 50 days.

  1. What would be the nominal cost of that credit?

Answer: The nominal cost of the credit is calculated as follows:

  • nominal cost of credit = percentage discount / (1 - percentage discount) * 365 / (full payment days - discount days)
  • nominal cost of credit = 3% / (1 - 3%) * 365 / (50 - 5)
  • nominal cost of credit = 25.09%

  1. What would be the effective cost of that credit?

Answer: Effective cost of trade credit = (1 + 3/97)365/45 – 1 = 1.2803 – 1 = 28.03%

  1. If the company could get the funds from a bank at a rate of 8%, interest paid monthly, based on a 365-day year, what would be the effective cost of the bank loan?

Answer: The effective annual rate of the bank loan is:

               (1+8%/12)12 – 1 = 8.30%

  1. Should Lancaster use bank debt or additional trade credit?

Answer: The effective cost of the bank loan is less than half the effective cost of the trade credit then the bank loan should be used.